The electric field at a distance of 2 m from an infinite line charge of charge density 4 nC/m is


1
Expert's answer
2022-03-20T18:48:52-0400

The electric field of linear charge

"E=\frac{\lambda}{2\pi\epsilon_0r}"

"E=\frac{4\times 10^{-9}}{2\times3.14\times8.85\times 10^{-12}\times 2}=36\:\rm N/C"